# Severity Controlled Text-to-Image Generative Model Bias Manipulation

### v1

@@ Abstract
Text-to-image (T2I) generative models are gaining wide popularity, especially in public domains. However, their intrinsic bias and potential malicious manipulations remain under-explored. Charting the susceptibility of T2I models to such manipulation, we first expose the new possibility of a dynamic and computationally efficient exploitation of model bias by targeting the embedded language models. By leveraging mathematical foundations of vector algebra, our technique enables a scalable and convenient control over the severity of output manipulation through model bias. As a by-product, this control also allows a form of precise prompt engineering to generate images which are generally implausible with regular text prompts. We also demonstrate a constructive application of our manipulation for balancing the frequency of generated classes - as in model debiasing. Our technique does not require training and is also framed as a backdoor attack with severity control using semantically-null text triggers in the prompts. With extensive analysis, we present interesting qualitative and quantitative results to expose potential manipulation possibilities for T2I models.
Index Terms:
Text-to-Image Models, Generative Models, Backdoor Attacks, Prompt Engineering, Bias
@@ 1 Introduction
The wide accessibility of large, multi-modal generative artificial intelligence (AI) models has led to a surge in interest surrounding AI by the general public, practitioners and researchers. Most sophisticated, state-of-the-art models are trained on large uncurated Internet data, which exposes them to harmful biases and representations publicized over time [1]. At the same time, public-facing applications are still bound by their black-box nature and their inherent bias characteristics are hard to quantify without conducting extensive experiments or gaining access to model parameters and training processes.
Due to its critical nature, generative model bias mitigation and manipulation is gaining instant attention of the research community [2, 3, 4, 5, 6, 7]. Model biasing/debiasing present opportunities in which inherent bias characteristics of a model may be leveraged to achieve specific goals. For instance, positive goals may include improving equity by focusing more on marginalised groups [2, 4, 5]. On the other hand, malicious backdoor manipulation is a more severe, negative objective of exploiting model bias [8, 3, 7].
@@
Multiple recent contributions emphasize that biased text-to-image (T2I) models lead to unfair social representations when generating images of people [9, 4, 10, 11, 12] as the generated content pertains to race, gender, cultural and geographic labels. This has severe negative implications when deployed in public applications, especially when marginalized groups are under-represented, or if harmful stereotypes are further emphasized by the model outputs. Concurrently, the stealthy and manipulative nature of backdoor attacks present an open challenge for the T2I generative domain [13, 14, 8, 7]. Such attacks can be used to propagate biases to an extreme degree, altering the target output upon the presence of stealthy triggers in the input. By considering the possibility of manipulating representations of people, objects, brands and ideologies, we can acknowledge the severe implications of backdoor attacks for T2I models. Generally, existing state-of-the-art bias manipulation and backdoor attack methods rely on training or fine-tuning the target models to manipulate their behaviour [2, 14, 8, 3, 4, 5, 10, 15, 7, 12]. This makes them less pragmatic. Moreover, the underlying techniques are generally computationally expensive and lack the flexibility of controlling the strength/severity of manipulation as this would mean re-training the model again and again.
@@
In this work, we expose the possibility of computationally efficient, dynamic bias manipulation that targets the embedded language model and allows tuning the severity of the bias manipulation using vector algebra for output interpolation and extrapolation - see Fig. 1. By shifting the language model embedding output using conventional vector algebra, our method is supported by solid mathematical foundations. Furthermore, the technique is scalable and applicable for generating precisely-engineered prompts. Thereby, it also enables generating images that would be otherwise implausible through regular language-based inputs. Guided by the consequences of bias exploitation in T2I models, we explore three impact perspectives supported by our bias manipulation method. (i)  Prompt engineering: tuning the language embedding space for a more precise image generation. (ii) Bias mitigation: shifting the language model embedding along $N$ vectors to balance class representations in the outputs of these models. (iii) Backdoor manipulation: exploiting the language model embedding space to change the perception of an object/class given the presence of a trigger in input.
@@
For our baseline prompt engineering and backdoor attack implementations, we use common class pairs based on CIFAR-10 dataset [16] labels, e.g., dog $\rightarrow$ cat, horse $\rightarrow$ deer. We also extend our prompt engineering experiments to consider manipulations as applied to gender, age and race classes. We carefully select our classes, noting that if a model provider chooses more sinister combinations, the consequences can be more severe. To improve social bias distributions we use social markers: gender (male and female), age (young and old) and race (white, black, asian). We define ‘labelled points’ in the language embedding space using discrete labels/classes. These points allow us to traverse from one point, e.g., dog, to another point, e.g., cat, in infinitesimally small intervals - something not plausible by relying just on natural language prompts. In summary, we contribute the following.
@@
    A technique to manipulate T2I models with language model embedding space interpolation (and extrapolation) for precise control over the embedding output. The manipulated embeddings allow generating images that would otherwise be implausible to describe with textual inputs.
@@
    Leveraging our manipulation approach, we contribute a method to shift the language embedding output to balance the frequency of classes being generated in the output using precise prompt engineering. Particularly, we demonstrate mitigating gender, age and racial biases through our method.
@@
    We extend our manipulation of the embedding outputs to implement a unique, dynamic, computationally-efficient backdoor attack that also enables severity tuning based on semantically-null triggers present in input prompts.
@@ 2 Background and Related Work
Generative Models: Text-to-image models leverage multimodal language and generative neural networks for user-guided, high-fidelity image synthesis. Seminal generative models were built on the foundation of solving the Nash equilibrium problem, i.e., learning an approximate probability distribution for some sample $\mathbf{x}$, ‘$\mathcal{P}_{model}(\mathbf{x})$’, that is close to the original probability distribution ‘$\mathcal{P}_{target}(\mathbf{x})$’ [17].
@@
Rombach et al. proposed the latent diffusion model [22], designing a high-fidelity T2I architecture while optimizing computational costs. This optimization results from separating the generative process into auto-encoding and diffusion sub-processes to learn low-dimensional latent representations and conceptual and semantic data compositions, respectively [22]. To improve flexibility and control, Brack et al. proposed a semantic guidance (SEGA) diffusion model in [19], exploiting multi-dimensional vector algebra to shift the generative process in the diffusion model space. We take a different approach and employ multi-dimensional vector algebra to manipulate the output of the language model (embedding) space. By targeting the language model output specifically, our bias manipulation methods are not necessarily limited to T2I downstream tasks.
@@
Stable diffusion, built from the latent diffusion model [22], leverages design methodologies/inspirations from DALL-E 2 and Imagen [23, 24] and has become one of the most popular T2I pipelines. Saharia et al. [24] proposed dynamic thresholding diffusion sampling technique and high guidance for high quality image generation, deploying their own variant of U-Net in the Imagen (and later Imagen 2) T2I models. Ramesh et al. and Betker et al. [23, 25], from OpentAI, proposed the DALL-E 2 and DALLE-3 hierarchical text-guided image generation models based on CLIP latents.
@@
A typical T2I pipeline contains: (a) a ‘tokenizer’ which converts input strings to tokenized representations, serving as input into a, (b) ‘text-encoder’ which projects the tokens onto an embedding space often using language-vision encoders. The text-embedding then serves as the conditional input to (c) a ‘generative model’, which through a text-guided, iterative latent deconstruction process, generates an image from an initial noisy representation. Many state-of-the-art diffusion models and T2I pipelines exploit the popular encoder-decoder U-Net architecture introduced in [26] for image synthesis. Our embedding manipulation occurs at the output of the language model as visualized in the high-level inference pipeline in Fig. 1.
@@
Bias and Backdoor Attacks: As the complexity and public awareness of machine learning and AI grows, so too does the discussion around bias and fairness in these domains [27, 9, 4, 10, 1]. Imbalanced social biases w.r.t to gender and race undoubtedly have a serious impact if not mitigated or at the very least quantified [10, 11, 4, 12]. In [9], Cho et al. proposed a method for evaluating social biases and visual reasoning of T2I models. Similarly, Naik et al. discussed social imbalances in T2I model outputs, focusing on race, age, gender and geographic markers [11]. Luccioni et al. proposed ‘StableBias’ for evaluating cultural and gender biases in T2I models [10]. Clemmer et al. proposed the instruction-following ‘PreciseDebias’ method, based on prompt engineering fundamentals and fine-tuning LLMs to mitigate demographic biases in T2I generative models [4]. Bolukbasi et al. explored the word embedding space and discussed the negative implications of male and female stereotypes and gender biases in [2] and used vector adjustments to reduce gender associations attached to an embedding [2]. Feng and Shah proposed an ‘epsilon-greedy’ re-ranking algorithm to mitigate gender biases and improve gender fairness in image searches [6].
@@
Backdoor attacks present an issue of extreme bias manipulation of target models and have been surveyed considerably across the literature [13, 28, 29]. While their applications in classification and decision-making systems has dominated, we find that the impacts of backdoor attacks on T2I models has recently gained traction [14, 8, 7, 15, 30, 31]. Chen et al. proposed a neural network Trojan attack on diffusion models “TrojDiff” in [14], exposing model vulnerabilities through an array of attacks to adjust the target model’s decision boundaries upon detection of an input trigger. Similarly, Chou et al. proposed a “BadDiffusion” backdoor attack that augments training and forward diffusion processes to adjust model output upon detection of an input trigger [8]. The Backdoor Attack on Generative Models (BAGM) method [7] targets T2I pipelines at various stages with three independent attacks to shift the bias towards popular brands [7]. The “Rickrolling the Artist” method details how CLIP-based text encoders in T2I models can be manipulated by exploiting rare trigger characters in the input prompt to manipulate the output [15].
Figure 2: Illustration of embedding space interpolation (and extrapolation) for (a) manual tuning of an input class towards a target class by traversing along the $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ vector as defined by Eq. (4), and (b) an arbitary example of the multi-cluster representation and how social biases can be balanced by traversing in $N$ directions in $\mathbb{E}^{n\times m}$ using Eq. (5). We choose seven clusters to balance their corresponding biases. From (a) to (b), we emphasize that our method is scalable, as evidenced by the number of clusters in each.
Figure 3: Visualizing how the embedding space can be exploited for a semantically-null, trigger-based backdoor attack. We show a representation of a semantically-null severity tuning dial within $\mathbb{E}^{n\times m}$ and assign severity values depending on the trigger token. Again, to implement these manipulations, an attacker does not require access to model weights and/or the network itself.
@@ 3.1 Definitions
Definition 3 - The embedding space: Given an $n\times m$-dimensional embedding space $\mathbb{E}^{n\times m}$, $\lambda(x,\phi_{L})$ outputs a text embedding $\mathbf{X}\in\mathbb{E}^{n\times m}$, where $\mathbf{X}=\{\mathbf{x}_{0},\mathbf{x}_{1},...,\mathbf{x}_{m}\}_{i}~{}\forall~{}i\in n$ i.e.,
$\mathbf{X}=\lambda(x,\phi_{L}).$ (2)
Definition 4 - Clustering in embedding space: In general, we exploit multiple clusters within $\mathbb{E}^{n\times m}$ to achieve various bias manipulation functions and embedding transformations. Given clusters $\mathbb{A}\in\mathbb{E}^{n\times m}$ and $\mathbb{B}\in\mathbb{E}^{n\times m}$, we define the corresponding cluster centroids $c_{i}$ as:
@@
Definition 5 - Precise prompt engineering: As visualized in Fig. 2(a), we can linearly control the bias manipulation severity using two pre-defined clusters $\mathbb{A}$ (trigger) and $\mathbb{B}$ (target) and their corresponding centroids $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$. For a language model output $\mathbf{X}\in\mathbb{A}$, we transform the output embedding such that:
@@
As visualized in Fig. 3, vector algebra is required to transform the embedding, given the semantically-null cluster centroid in $\mathbb{E}^{n\times m}$. So, given a text embedding $\mathbf{X}\in\mathbb{E}^{n\times m}$ and the semantically-null trigger and target cluster centroids - $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ respectively, we can derive $\overline{\mathbf{X}}$ as:
@@
3.2 Precise Prompt Engineering through Embedding Interpolation
As illustrated in Fig. 2(a), on a foundational level, we can interpolate the language model embedding outputs using two clusters $\mathbb{A}$ and $\mathbb{B}$. Given the high-dimensionality of the embedding space and the complexity of natural language, these labelled points can have a varied range in terms of similarity, e.g., dogs vs aliens or two different breeds of the same species.
As modelled by Eq. (4) and visualized in Fig. 2(a), by defining $c_{\mathbb{A}},c_{\mathbb{B}}\in\mathbb{E}^{n\times m}$, we can engineer the prompt and interpolate the output along the $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ vector, using some scalar $\mathcal{S}$. Equation (4) does not imply that $\mathcal{S}$ is bound by the distance between the two centroids. In fact, by defining $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$, we can extrapolate beyond $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ in forward ($\mathcal{S}>1$) and reverse ($\mathcal{S}<0$) directions. Effectively, through these vector transformations, we are controlling model biases without training or fine-tuning the network, purposefully shifting it towards or away from some target class $\mathbb{B}$ using $\mathcal{S}$. From an application perspective, this method has positive and negative implications depending on the intent.
This allows us to precisely engineer the prompt via the language model output, such that the images generated by a downstream diffusion model for example, could be representative of something that is near impossible (or extremely difficult) to define using a purely textual input. Through our prompt engineering approach, we can effectively create continuous spectra using binary points $c_{\mathbb{A}},c_{\mathbb{B}}\in\mathbb{E}^{n\times m}$. Then, by increasing the number of known cluster centroid locations in $\mathbb{E}^{n\times m}$, we are granted more flexibility and control over the embedding space and thus, over T2I model biases.

### v2

@@ Abstract
Text-to-image (T2I) generative models have gained increased popularity in the public domain. While boasting impressive user-guided generative abilities, their black-box nature exposes users to intentionally- and intrinsically-biased outputs. Bias manipulation (and mitigation) techniques typically rely on careful tuning of learning parameters and training data to adjust decision boundaries to influence model bias characteristics, which is often computationally demanding. We propose a dynamic and computationally efficient manipulation of T2I model biases by exploiting their rich language embedding spaces without model retraining. We show that leveraging foundational vector algebra allows for a convenient control over language model embeddings to shift T2I model outputs and control the distribution of generated classes. As a by-product, this control serves as a form of precise prompt engineering to generate images which are generally implausible using regular text prompts. We demonstrate a constructive application of our technique by balancing the frequency of social classes in generated images, effectively balancing class distributions across three social bias dimensions. We also highlight a negative implication of bias manipulation by framing our method as a backdoor attack with severity control using semantically-null input triggers, reporting up to 100% attack success rate.
Keywords:
Text-to-Image Models Generative Models Bias Prompt Engineering Backdoor Attacks
@@ 1 Introduction
The accessibility of large, multi-modal generative artificial intelligence (AI) models has led to a widespread surge in AI interest. Most sophisticated, state-of-the-art models are trained on large uncurated Internet data, which exposes them to harmful biases and representations publicized over time [24]. At the same time, public-facing applications are still bound by their black-box nature and their biases are hard to quantify without conducting extensive experiments or gaining access to model and training parameters. Due to its critical nature, generative model bias mitigation and manipulation is gaining instant attention of the research community [4, 9, 10, 11, 12, 35]. Model biasing/debiasing present opportunities in which inherent bias characteristics of a model may be leveraged to achieve specific goals. Positive goals may include improving equity by focusing more on marginalised groups [4, 10, 11]. On the other hand, malicious backdoor attacks are a more severe, negative objective of manipulating model bias [8, 9, 35].
@@
Recent contributions emphasize that biased text-to-image (T2I) models lead to unfair social representations when generating images of people [7, 10, 22, 25, 34] as the generated content pertains to race, gender, cultural and geographic labels. This has severe negative implications for public-facing applications, especially when marginalized groups are under-represented, or if harmful stereotypes are further emphasized through model outputs. Concurrently, the stealthy and manipulative nature of backdoor attacks present an open challenge for the T2I generative domain [1, 6, 8, 35]. Such attacks can be used to propagate biases to an extreme degree, altering the target output upon the presence of stealthy triggers in the input. By considering the possibility of manipulating representations of people, objects, brands and ideologies, we can acknowledge the severe implications posed by backdoor attacks on T2I models. Generally, existing state-of-the-art bias manipulation and backdoor attack methods rely on training or fine-tuning the target models to manipulate their behaviour [4, 6, 8, 9, 10, 11, 22, 33, 35, 34]. This makes them less pragmatic. Moreover, the underlying techniques are generally computationally expensive and lack the flexibility of controlling the strength/severity of manipulation as this would require continuous re-training of the model.
@@
Multidimensional interpolation has been used in computer graphics and more recently in T2I and diffusion model literature [32, 14, 18, 36]. Our method exploits language model embedding vectors and traditional vector algebra to manipulate T2I model biases in positive and negative contexts. Thus, we expose the possibility of a dynamic, computationally efficient bias manipulation strategy as visualized in Fig. 1. The technique is scalable and applicable for generating precisely-engineered prompts. Thereby, it also enables generating images that would be otherwise implausible through regular language-based inputs. Guided by the consequences of bias exploitation in T2I models, we explore three impact perspectives supported by our bias manipulation method. (i)  Prompt engineering: Measuring the fundamental abilities of our manipulation method to exploit the language embedding space for controlled, precise image generation and class manipulation. (ii) Bias mitigation: shifting the language model embedding along $N$ vectors to balance class representations in model outputs. (iii) Backdoor manipulation: exploiting the embedding space to intentionally shift the perception (bias) of an object/class when a semantically-null trigger is present in the input.
@@
For our fundamental prompt engineering and backdoor attack implementations, we use common class pairs based on CIFAR-10 dataset [17] labels, e.g., dog $\rightarrow$ cat, horse $\rightarrow$ deer. These class combinations are benign by choice. If a model provider chooses more sinister combinations, the consequences would be more severe. To mitigate social biases we use social markers: gender (male and female), age (young and old) and race (white, black, asian). We define ‘labelled points’ in the language embedding space using discrete labels/classes. These points allow us to traverse from one point, e.g., dog, to another point, e.g., cat, in infinitesimally small intervals. In summary, we contribute the following.
@@
    A language model embedding output manipulation technique exploiting vector transformations. Cluster-defined labelled points in the embedding space allow for precise control over embeddings and thus, the generated image.
@@
    Leveraging our manipulation technique, we contribute a method to shift the language embedding output to balance the frequency of classes in generated classes. Particularly, we demonstrate mitigating gender, age and racial biases.
@@
    We extend our manipulation of the embedding outputs to implement a unique, dynamic, computationally-efficient backdoor attack that also enables severity tuning based on semantically-null triggers present in input prompts. This highlights the negative implications of controllable bias manipulations.
@@ 2 Background and Related Work
Generative Models Text-to-image models leverage multimodal language and generative neural networks for user-guided, high-fidelity image synthesis. Seminal generative models were built on the foundation of solving the Nash equilibrium problem, i.e., learning an approximate probability distribution for some sample $\mathbf{x}$, ‘$\mathcal{P}_{model}(\mathbf{x})$’, that is close to the original probability distribution ‘$\mathcal{P}_{target}(\mathbf{x})$’ [13].
@@
Rombach et al. proposed the high-fidelity latent diffusion model architecture with optimized computational costs [28]. This optimization results from separating the generative process into auto-encoding and diffusion sub-processes to learn low-dimensional latent representations and conceptual and semantic data compositions, respectively [28]. Stable diffusion, built from the latent diffusion model [28], leverages design methodologies/inspirations from DALL-E 2 and Imagen [27, 30]. Saharia et al. [30] proposed dynamic thresholding diffusion sampling technique and high guidance for high quality image generation, deploying their own variant of U-Net in the Imagen (and later Imagen 2) T2I models. Ramesh et al. and Betker et al. present DALL-E 2 (and DALL-E 3) models, a high-fidelity, hierarchical text-guided image generation models based on CLIP latents [27, 3].
@@
A typical T2I pipeline contains: (a) a ‘tokenizer’ which converts input strings to tokenized representations, serving as input into a, (b) ‘text-encoder’ which projects the tokens onto an embedding space often using language-vision encoders like the CLIP model [26]. The text-embedding then serves as the conditional input to (c) a ‘generative model’, which through a text-guided, iterative latent deconstruction process, generates an image from a random initial noisy representation. Many state-of-the-art diffusion models and T2I pipelines exploit the popular encoder-decoder U-Net architecture introduced in [29] for image synthesis. Our embedding manipulation occurs at the output of the language model as visualized in the high-level inference pipeline in Fig. 1.
Spatial Interpolation: Recently, spatial interpolation has been deployed to walk the latent space in diffusion models and to grant users improved generative control. Brack et al. proposed a semantic guidance (SEGA) diffusion model in [5], exploiting multi-dimensional vector algebra to shift the generative process in the diffusion model space. Guo et al. propose Smooth Diffusion [14] to improve the editing potential of T2I generative models. They re-train a stable diffusion U-Net while freezing the text-encoder, optimizing the ratio between the input latent and the variation in the output prediction [14]. Kwon et al. demonstrate the presence of semantic latent spaces in frozen diffusion models, deploying an asymmetric reverse process to modify generated images in the latent space during inference [18]. Our approach differs to [18] as we only manipulate the embedding output, leaving the diffusion process untouched. Wang et al. propose generating interpolated images from the diffusion model latent space using two real image points as vector extremes [36]. While [36] proposes translations between real images, their method could feasibly be extended to generated images given a consistent latent space. Across [14, 18, 36], we observe that unlike our method, bias manipulations were not a key consideration. Furthermore, [14, 18, 36] evidence that the latent space is more commonly used for manipulating generated images, whereas we focus on the embedding space.
@@
Bias and Backdoor Attacks: As the complexity and public awareness of machine learning and AI grows, so too does the discussion around bias and fairness in these domains [2, 7, 10, 22, 24]. Imbalanced social biases w.r.t gender and race undoubtedly have a serious impact if not mitigated or at the very least quantified [22, 25, 10, 34, 39]. In [7], Cho et al. proposed a method for evaluating social biases and visual reasoning of T2I models. Similarly, Naik et al. discussed social imbalances in T2I model outputs, focusing on race, age, gender and geographic markers [25]. Luccioni et al. proposed ‘StableBias’ for evaluating cultural and gender biases in T2I models [22]. Clemmer et al. proposed the instruction-following ‘PreciseDebias’ method, based on prompt engineering fundamentals and fine-tuning LLMs to mitigate demographic biases in T2I generative models [10]. Bolukbasi et al. explored the word embedding space and discussed the negative implications of male and female stereotypes and gender biases in [4] and used vector adjustments to reduce gender associations attached to an embedding [4]. Feng and Shah proposed an ‘epsilon-greedy’ re-ranking algorithm to mitigate gender biases and improve gender fairness in image searches [12]. Zhang et al. propose the ITI-GEN method, using reference images to guide prompt learning without fine-tuning the text-to-image model [39]. Our method does not require reference images, relying instead on prompt-based clusters that already exist in embedding spaces of pre-trained models.
@@
Backdoor attacks present an issue of extreme bias manipulation of target models and have been surveyed considerably across the literature [1, 16, 20], with backdoor attacks on T2I models having recently gained significant traction [6, 8, 35, 33, 38, 40]. Chen et al. proposed a neural network Trojan attack on diffusion models “TrojDiff” in [6], exposing model vulnerabilities through an array of attacks to adjust the target model’s decision boundaries upon detection of an input trigger. Similarly, Chou et al. proposed a neural network backdoor attack “BadDiffusion” that augments training and forward diffusion processes to adjust model output upon detection of an input trigger [8]. The “BAGM” method [35] targets T2I pipelines at various stages with three independent attacks to shift the bias towards popular brands [35]. Struppek et al. propose exploiting rare triggers in CLIP-based text encoders to influence T2I model outputs [33].
@@ 3.1 Definitions
Definition 3 - The embedding space: Given an $n\times m$-dimensional embedding space $\mathbb{E}^{n\times m}$, the language model ‘$\lambda(x,\phi_{L})$’ outputs a text embedding $\mathbf{X}\in\mathbb{E}^{n\times m}$, where $\mathbf{X}=\{\mathbf{x}_{0},\mathbf{x}_{1},...,\mathbf{x}_{m}\}_{i}~{}\forall~{}i\in n$.
Definition 4 - Embedding clusters: In general, we exploit multiple clusters within $\mathbb{E}^{n\times m}$ to achieve various bias manipulation functions and embedding transformations. Given clusters $\mathbb{A}\in\mathbb{E}^{n\times m}$ and $\mathbb{B}\in\mathbb{E}^{n\times m}$, we define the corresponding cluster centroids $c_{i}$ as:
@@
Figure 2: Illustration of embedding manipulations for (a) manual tuning of an input class towards a target class by traversing along the $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ vector as defined by Eq. (4), and (b) an arbitrary example of the multi-cluster representation and how social biases can be balanced by traversing along $N$ directions in $\mathbb{E}^{n\times m}$ using Eq. (5).
Definition 5 - Embedding transformations: As visualized in Fig. 2(a), we can linearly control the bias manipulation severity using two pre-defined clusters $\mathbb{A}$ (trigger) and $\mathbb{B}$ (target) and their corresponding centroids $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$. For a language model output $\mathbf{X}\in\mathbb{A}$, we transform the output embedding such that:
@@
As visualized in Fig. 3, vector algebra is required to transform the embedding, given the additional semantically-null cluster centroid in $\mathbb{E}^{n\times m}$. So, given a text embedding $\mathbf{X}\in\mathbb{E}^{n\times m}$ and the semantically-null trigger and target cluster centroids - $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ respectively, we can derive $\overline{\mathbf{X}}$ as:
@@
Figure 3: Visualizing how the embedding space can be exploited for a semantically-null, trigger-based backdoor attack. We show a representation of a semantically-null severity tuning dial within $\mathbb{E}^{n\times m}$ and assign severity values depending on the trigger token.
3.2 Prompt Engineering Through Embedding Manipulations
As illustrated in Fig. 2(a), we can apply vector transforms to the embedding output using two clusters $\mathbb{A}$ and $\mathbb{B}$. Given the high-dimensionality of the embedding space and the complexity of natural language, the semantic relationship between labelled points can vary greatly.
As modelled by Eq. (3) and visualized in Fig. 2(a), by defining $c_{\mathbb{A}},c_{\mathbb{B}}\in\mathbb{E}^{n\times m}$, we can modify the embedding and interpolate the output along the $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$ vector, using some scalar $\mathcal{S}$. Equation (3) does not imply that $\mathcal{S}$ is bound by the distance between the two centroids. In fact, by defining $\overrightarrow{c_{\mathbb{A}}c_{\mathbb{B}}}$, we can extrapolate beyond $c_{\mathbb{A}}$ and $c_{\mathbb{B}}$ in forward ($\mathcal{S}>1$) and reverse ($\mathcal{S}<0$) directions. Effectively, through these vector transformations, we are controlling model biases without training or fine-tuning the network, purposefully shifting outputs towards or away from some target class $\mathbb{B}$ using $\mathcal{S}$. From an application perspective, this method has positive and negative implications depending on intent.
Our language model embedding manipulation allows us to generate images that may be near-impossible to define using purely textual prompts. By identifying labelled points in the embedding space, we can effectively convert any binary attribute to a continuous one. By increasing the number of known cluster centroid locations in $\mathbb{E}^{n\times m}$, we are granted more flexibility and control over the embedding space and thus, over T2I model biases and generated content.